Can someone please show me how to graph this,
x=-3, y=200
x=0, y=100
x=3, y=59
x=6, y=25
x=9, y=12.5
x=12, y=6.25
x=-6, y=450

Answer by Alan3354(69443)

Plot the points, draw a curve thru them.
====================
It appears to be a parabola.
To find the equation:
y = ax^2 + bx + c
----
Use any 3 points.
100 = a*0 + b*0 + c --> c = 100
---
59 = a*9 + b*3 + 100
25 = a*36 + b*6 + 100
----
Solve for a & b
9a + 3b = -41
36a + 6b = -75
-----
36a + 12b = -164 --- 1st eqn times 4
36a + 6b = -75
-------------------------- Subtract
6b = -89
b = -89/6
---
a = 7/18
-----------------
y = 7x^2/18 - 89x/6 + 100
=====================
That doesn't fit all the points, only the 3 used --> not a parabola